### v1

# DynOPETs: A Versatile Benchmark for Dynamic Object Pose Estimation and Tracking in Moving Camera Scenarios

Abstract
In the realm of object pose estimation, scenarios involving both dynamic objects and moving cameras are prevalent. However, the scarcity of corresponding real-world datasets significantly hinders the development and evaluation of robust pose estimation models. This is largely attributed to the inherent challenges in accurately annotating object poses in dynamic scenes captured by moving cameras. To bridge this gap, this paper presents a novel dataset DynOPETs and a dedicated data acquisition and annotation pipeline tailored for object pose estimation and tracking in such unconstrained environments. Our efficient annotation method innovatively integrates pose estimation and pose tracking techniques to generate pseudo-labels, which are subsequently refined through pose graph optimization. The resulting dataset offers accurate pose annotations for dynamic objects observed from moving cameras. To validate the effectiveness and value of our dataset, we perform comprehensive evaluations using 18 state-of-the-art methods, demonstrating its potential to accelerate research in this challenging domain. The dataset will be made publicly available to facilitate further exploration and advancement in the field.
I INTRODUCTION
Object pose estimation, which involves determining an object’s 6-DoF pose (position and orientation) in 3D space, is a fundamental research challenge in computer vision and robotics. The task is crucial for various applications in embodied intelligence, augmented reality (AR), and mixed reality (MR) systems [1], such as robotic manipulation [2], dexterous hand control, and human-object interaction [3]. Despite significant advancements in recent years driven by the emergence of benchmark datasets, existing real-world datasets exhibit notable limitations. Most of the existing datasets predominantly assume static conditions - either fixed camera positions or stationary objects. This assumption contrasts sharply with real-world requirements where simultaneous camera ego-motion and object dynamics must be addressed.
The limited availability of dynamic pose estimation datasets stems primarily from annotation challenges inherent to moving-camera-dynamic-object configurations. While static-object scenarios allow for indirect pose annotation via the state-of-the-art camera pose estimation technique, such as SLAM algorithm [4] or fiducial marker-based localization method [5], dynamic settings require direct object motion tracking. In such dynamic settings, existing solutions for pose annotation rely on either labor-intensive manual annotation [6] or expensive motion capture (Mocap) systems in controlled environments [7].
Figure 1: The full set of dynamic objects in our dataset.
Another approach involves using state-of-the-art object pose estimation methods to generate high-quality pseudo-labels for annotation. However, many common objects in daily life, such as those with symmetric geometries, transparent or reflective materials, or sparse/repetitive patterns, pose significant challenges for those algorithms. As a result, a significant fraction of the object pose annotations still require manual refinement due to the low quality of the pseudo-labels.
To address the issue of scarce real-world datasets in this challenging scenario, we propose a novel dataset that includes diverse object categories and dynamic scenes. The dataset primarily consists of RGB-D video sequences, object CAD models, and high-quality object and camera pose annotations, supporting various tasks in the field of object pose estimation. To tackle the challenges of pose annotation, we introduce a novel automated pose annotation pipeline leveraging the fusion of absolute and relative object pose estimates, significantly improving data labeling efficiency and reducing human effort. Finally, we systematically investigate existing state-of-the-art methods and comprehensively benchmark their performance in dynamic scenarios, demonstrating the value of our work. We believe that our proposed dataset and efficient pose annotation pipeline can fill the gap in dynamic object pose estimation datasets, thereby promoting the development of applications in the fields of embodied intelligence and AR/MR. Our main contributions are summarized as follows:
  • We propose an efficient pose annotation approach that integrates an absolute pose estimator with a global EKF smoother, a relative pose estimator based on point tracking, and an object pose graph optimization module. This pipeline enables accurate, marker-free dynamic object pose labeling, without extensive manual effort.
  • We present DynOPETs, a comprehensive object pose estimation dataset comprising RGB-D sequences of 175 distinct object instances undergoing simultaneous camera and object motion, with synchronized 6-DoF pose annotations for both camera and object trajectories.
  • We conduct a systematic benchmark of various taxonomy-specific pose estimation and object pose tracking methods tailored to the unique characteristics of our dataset.
TABLE I: Comparison between existing 17 datasets with our proposed dataset. The table summarizes key attributes used for comparing datasets: object categories, whether the data is synthetic or real-world, data modalities (RGB, Depth), object details (number of instances, presence of dynamic objects, presence of moving cameras, CAD model availability), whether the videos are marker-free, and dataset size (number of videos, images, and annotations).
Dataset Cat. Real Modality Object Details Marker-Free Vid. Img. Anno.
RGB Depth Num. CAD Dyna-Obj Mov-Cam
CAMERA25 [8] 6 1085 - 300K 4M
Omni6D [9] 166 4688 - 0.8M 5.6M
SOPE [10] 149 4162 - 475K 5M
YCB-Video [11] - 21 92 133K 0.6M
T-LESS [12] - 30 20 48K 48K
Linemod [13] - 15 - 18K 15k
StereoOBJ-1M [14] - 18 182 393K 1.5M
YCBInEOAT [6] - 5 9 7449 3.7K
HOT3D [7] - 33 425 3.7M 6M
UOPE-56 (Ours) - 56 56 64K 180K
REAL275 [8] 6 42 18 8K 35K
Wild6D [15] 5 162 486 10K 10K
PhoCaL [16] 8 60 24 3.9K 91K
HANDAL [17] 17 212 2.2K 308K 1.2M
HouseCat6D [18] 10 194 41 23.5K 160K
ROPE [10] 149 581 363 332K 1.5M
COPE-119 (Ours) 6 119 119 140K 350K
II RELATED WORK
According to  [19], We categorize object pose estimation methods into 4 classes: instance-level object pose estimation (IOPE), unseen object pose estimation (UOPE), category-level object pose estimation (COPE), and object pose tracking (OPT). This section discusses representative and state-of-the-art methods within each class, followed by an overview of recently published datasets relevant to these areas.
II-A Object Pose Estimation and Tracking Methods
II-A1 Instance-level Object Pose Estimation
These methods predict the pose of only those objects present in the training set, often relying on priors such as CAD models or posed reference views during inference [20]. Object pose estimation can be achieved through template matching [21] or feature correspondence matching [22] given these priors. Some approaches directly predict object pose using neural networks [20], while others employ voting schemes [23]. Instance-level methods are the most mature but also the most restrictive due to their inability to generalize to novel objects. Consequently, research has shifted towards the more generalizable UOPE and COPE methods, which are the primary focus of this paper and our proposed dataset.
II-A2 Unseen Object Pose Estimation
Similar to IOPE methods, UOPE methods rely on CAD models or posed reference views as priors. However, UOPE methods can generalize to novel objects, enabling pose prediction for unseen objects during inference. Recent works in this area are broadly categorized into two main strategies: feature matching and template matching-based methods. Feature matching methods [24, 25, 26, 27] extract robust image features and establish correspondences with generic, object-agnostic CAD model features. These approaches enable zero-shot estimation but may rely heavily on feature discriminability. In contrast, template matching methods [28, 29, 30, 1] leverage extensive sets of object templates to achieve strong generalization across novel objects.
II-A3 Category-level Object Pose Estimation
COPE methods typically handle novel objects within the same category by learning category-specific shape representations, eliminating the reliance on CAD models and reference views. Wang et al. [8] pioneered this approach by introducing the normalized object coordinate space (NOCS), a canonical representation where all objects within the same category are consistently aligned in terms of size and orientation. Aligning predicted canonical representations with depth points allows for the estimation of object pose.
Recent advancements have explored different network designs, such as using vision transformers [31] or diffusion models [32] to predict better canonical representations. Several works [33, 34, 35, 32, 36] employ deformed categorical shape templates to improve canonical representations, while others [37, 10] employ diffusion models to formulate COPE as a conditional generative task. Furthermore, some methods directly predict object pose, employing either enhanced geometric feature extractors [38] or a combination of geometric and semantic priors [39, 40, 41].
II-A4 Object Pose Tracking
Unlike single-frame object pose estimation methods, object pose tracking leverage temporal information across video sequences to estimate object poses [19]. Typically, tracking methods also require posed reference views or CAD models. While classical methods [42] often track object poses using geometric or feature-based approaches, recent methods [43] jointly track and refine object shape through neural implicit representations, providing robust performance. Moreover, Wen et al. [1] leverage large-scale pre-trained models and efficiently track unseen objects using large-scale synthetic training data for enhanced generalization.
II-B Object Pose Estimation Datasets
Object pose estimation datasets have evolved significantly over the years. We show a dataset comparison in table I. The following is a brief overview.
Linemod [13] introduced a standard instance-level RGB-D benchmark featuring texture-less objects with limited occlusions in relatively simple scenarios. YCB-Video [11] subsequently expanded realism with household objects. T-Less [12] introduced texture-less but symmetric objects in challenging industrial scenarios, highlighting pose ambiguity issues. REAL275 [8] established the first real-world benchmark for category-level pose estimation by introducing canonical object representations across 6 categories. StereOBJ-1M [14] further tackled challenging materials with a much larger scale, such as transparent and reflective objects. To further diversify modalities, PhoCaL [16] provided multimodal images focusing on challenging transparent and reflective household items with 8 categories. Wild6D [15] advanced realism further by capturing large-scale videos of objects in the wild. More recently, HANDAL [17] focused on graspable objects, collecting 17 categories of manipulable items in cluttered scenes via multi-view imagery and annotating poses with [43]. HOT3D [3] was the first to address egocentric scenarios featuring dense hand–object interactions. However, it incorporates a limited variety of objects and constrained scene settings. HouseCat6D [18] delivered highly accurate annotations for 10 categories of household objects using polarization imagery to enhance robustness to reflections. OmniNOCS [31] was a unified NOCS dataset that integrated data from multiple domains and featured over 90 object classes, making it the largest NOCS dataset to date. Omni6D [9] and ROPE [10] also provided a large-scale dataset for object pose estimation.
Figure 2: Overview. The left block is CAD model scanning and preprocessing, as well as hand-eye calibration. The right block is the data acquisition and efficient object pose annotation pipeline. The dynamic object is captured by a moving RGB-D camera, whose ground truth pose annotation is calculated by fusing refined absolute object pose estimates and relative object pose estimates. The camera pose is provided by a Mocap.
III Data Acquisition
III-A Overview
The proposed DynOPETs consists of 175 RGB-D video sequences of distinct everyday objects captured on typical desktop backgrounds with sufficient visual features. Each sequence comprises roughly 500 frames recorded at 15 Hz, totaling approximately 210K RGB-D frames. Additionally, we provide over 530K annotations, included instance segmentations, camera intrinsic and extrinsic parameters, object poses, and camera poses. The dataset is divided into two subsets. The first, COPE-119, comprises 119 sequences of objects belonging to 6 categories commonly found in the COPE benchmark: bottles, bowls, cameras, cans, laptops, and mugs. The second, UOPE-56, contains 56 sequences of other common household objects (illustrated in  fig. 1) and is designed for evaluating the performance of UOPE algorithms.
III-B Hardware Setup and Objects
RGB-D data was captured using a consumer-grade iPad Pro with the hardware-synchronized Structure Sensor Pro [44] depth camera. CAD models for all 175 objects were scanned using the sensors’ integrated RGB-D 3D modelling software. Following mesh reconstruction, objects were manually aligned so that their centers coincided with the origin of the 3D coordinate system, and their front, top, and side views aligned with the x, y, and z axes, respectively. Consistent with the definition in Wang et al. [8], the coordinate system of an object’s CAD model is defined as Object Local Coordinates (OLC) system, also known as the canonical space in the COPE domain, as illustrated in Figure 3. Additionally, we employed SAM2 [45] to automatically segment objects in these frames.
III-C Sensor Calibration and Synchronization
Intrinsic and extrinsic camera parameters were calibrated using the depth camera SDK, which provides simultaneous calibration of both RGB and depth cameras. We developed a custom camera driver based on the SDK that warps the depth map to the RGB camera’s coordinate frame, resulting in co-registered RGB-D video sequences. A Mocap system provided camera poses. The relative transformation between the Mocap marker frame and the camera’s optical center was calibrated using a hand-eye calibration procedure, yielding camera poses within the Mocap coordinate system. A temporal offset existed between the Mocap system’s timestamps and the RGB-D frame timestamps. To resolve this, we used visual odometry (VO) [4] to estimate the camera pose for each frame and registered these VO-derived poses to the Mocap poses. This registration process established a mapping between each camera frame and the corresponding Mocap timestamp.
III-D Efficient Object Pose Annotation
To provide context for our approach, we begin with key definitions. Firstly, the absolute object pose, which we aim to annotate, is defined as the Sim(3) transformation (scale, rotation, and translation) required to align an object from its OLC system to the camera frame. Since our CAD models and corresponding point clouds are scaled to real-world dimensions, this Sim(3) simplifies to an SE(3) transformation (rotation and translation). Secondly, the relative object pose describes the SE(3) transformation between the object’s poses at two different time steps in the camera coordinate system. These two definitions are directly relevant to the following two approaches for obtaining the object pose annotations.
III-D1 Absolute Object Pose Estimation and Refinement
The first approach involves directly generating high-quality pseudo pose annotations using existing state-of-the-art algorithms. Model-based method FoundationPose [1] stands out in this category, leveraging CAD models to output object motion. Although generally accurate, FoundationPose [1] can occasionally generate discontinuous pose estimates. To address this issue, we adopt a global EKF [46] with Rauch-Tung-Striebel (RTS) smoother [47] to refine the object pose predictions, mitigating pose jitters. The RTS smoother employs a constant velocity model for both translation $p\in\mathbf{R}^{3}$ and rotational quaternion $q\in\mathbb{S}^{3}$, where the state variables are defined as $\mathbf{x}=[p,q,\upsilon,\omega]$, where $\upsilon\in\mathbb{R}^{3},\omega\in\mathbb{R}^{3}$ refers the velocity and angular velocity. We first perform a forward pass eq. 2 using the standard EKF prediction and update equations to obtain filtered state estimates $\mathbf{x}_{t}$ from step $t-1$ to step $t$, then, a backward pass eq. 3 using the RTS smoother refines these estimates from step $t+1$ to $t$:
$\hat{\mathbf{x}}_{t}=\mathbf{F}\mathbf{x}_{t-1},$ (1)
$\mathbf{x}_{t}=\hat{\mathbf{x}}_{t}+\mathbf{K}_{t}\,(\mathbf{z}_{t}-\mathbf{H}\hat{\mathbf{x}}_{t}),$ (2)
$\mathbf{x}_{t}=\mathbf{x}_{t}+\mathbf{G}_{t}(\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1}),$ (3)
$\mathbf{G}_{t}=\mathbf{P}_{t}\mathbf{F}^{T}\mathbf{P}_{t+1}^{-1},$ (4)
where $\mathbf{z}$ refers to the observed pose, and $\mathbf{F}$ and $\mathbf{H}$ represent the state transition and observation matrices, respectively. $K$ is the extended Kalman gain used in the update step [46], $\mathbf{P}$ denotes the covariance matrix, and $\mathbf{G}$ represents the smoothing gain. After our refinement, the pose annotations of objects with rich textures and distinctive shapes are now accurate and smooth, while for those objects exhibiting significant symmetry or lacking distinctive texture, such as objects within the bowl category, the pose annotation is still not high-quality enough.
III-D2 Relative Object Pose Estimation via Point Tracking
To address the challenges posed by these objects, we explore a second approach that involves first estimating the absolute object pose in the initial frame and then iteratively computing the frame-to-frame relative object pose to derive the pose annotation for the entire sequence.
BundleTrack [42] serves as a representative example of this approach, employing keypoint detection and matching to establish 2D correspondences. These 2D correspondences are then back-projected into 3D space using sensor depths, forming 3D-3D correspondences that enable the computation of relative object poses via point registration techniques. Our method adopts a similar strategy but replaces the keypoint detection and matching component with a state-of-the-art Track Any Point (TAP) model [48]. TAP models accept a video sequence as input and output 2D correspondence points for a group of specified query points in the reference frame across all frames. Compared to traditional feature extraction and matching methods, TAP model [48] is better equipped to leverage the temporal information inherent in video sequences, enabling them to effectively handle objects with limited texture. While this method produces smoother object pose estimates and is more robust to symmetry ambiguities, it is susceptible to cumulative errors, leading to drift in the estimated trajectory over time.
III-D3 Object Pose Graph Optimization
Recognizing the complementary nature of the two annotation methods discussed before, we construct a pose graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ to integrate them into a unified framework. The graph consists of nodes $\mathcal{V}$ and edges $\mathcal{E}$(illustrated in fig. 3). Each node $i\in\mathcal{V}$ represents the pose of an object at a specific time step, denoted as $\mathbf{T}_{i}\in SE(3)$, where $\mathbf{T}_{i}$ represents the transformation from the object’s local coordinate system to the camera coordinate system at time $i$. In essence, objects in OLC and their corresponding camera frames at each time step become the nodes in this graph.
The edges $\mathcal{E}$ represent constraints between these object poses. We utilize two types of edges: absolute pose edges and relative pose edges. Absolute pose edges, denoted as $(\text{OLC},i)\in\mathcal{E}$, are derived from FoundationPose refined by the EKF/RTS smoother. Relative pose edges, denoted as $(i,j)\in\mathcal{E}$, connect two nodes representing the same object at different time steps. These are derived from the point tracking method (TAP). We use $\mathbf{z}_{\text{OLC},i},\mathbf{z}_{i,j}\in SE(3)$ to represent the associated measurements of these edges, and $\mathbf{\Omega}_{i,j}$ and $\mathbf{\Omega}_{\text{OLC},i}$ for the corresponding information matrices. Note that in our implementation, $\mathbf{\Omega}_{\text{OLC},i}$ is constant and $\mathbf{\Omega}_{i,j}$ are calculated from the Jacobian of the point set registration. Also, absolute pose edges with poor measurements can be easily removed manually.
By encoding both absolute and relative poses as edges connecting these nodes, we form a graph with closed loops. Then we need to finding the set of object poses $\mathbf{T}=\{\mathbf{T}_{i}\}_{i\in\mathcal{V}}$ that minimizes the following cost function:
$\mathcal{F}(\mathbf{T})$ $=\sum_{(\text{OLC},i)\in\mathcal{E}}\mathbf{r}_{\text{OLC},i}^{T}(\mathbf{T})\mathbf{\Omega}_{\text{OLC},i}\mathbf{r}_{\text{OLC},i}(\mathbf{T})$ (5)
$+\sum_{(i,j)\in\mathcal{E}}\mathbf{r}_{i,j}^{T}(\mathbf{T})\mathbf{\Omega}_{i,j}\mathbf{r}_{i,j}(\mathbf{T}),$ (6)
where $\mathbf{r}_{\text{OLC},i}(\mathbf{T})$ and $\mathbf{r}_{i,j}(\mathbf{T})$ are the residual error terms for absolute and relative pose edges, respectively, defined as:
$\mathbf{r}_{\text{OLC},i}(\mathbf{T})=\log(\mathbf{z}_{\text{OLC},i}^{-1}\mathbf{T}_{i}),$ (7)
$\mathbf{r}_{i,j}(\mathbf{T})=\log(\mathbf{z}_{i,j}^{-1}\mathbf{T}_{i}^{-1}\mathbf{T}_{j}),$ (8)
Here, $\log(\cdot)$ denotes the matrix logarithm, mapping an $SE(3)$ transformation to its Lie algebra representation.
The optimization problem, therefore, is to find:
$\mathbf{T}^{*}=\arg\min_{\mathbf{T}}\mathcal{F}(\mathbf{T}).$ (9)
This non-linear least squares problem can be solved using Gauss-Newton or Levenberg-Marquardt algorithms. By minimizing the cost function, we obtain refined object poses that are globally consistent, leveraging the strengths of both absolute and relative pose estimation methods. The manual effort required by our method is primarily limited to identifying and removing inaccurate absolute pose estimates, which minimizes human intervention and promotes annotation efficiency.
Figure 3: Pose graph Optimization. The orange arrows represent absolute pose edges $(\text{OLC},i)$ and the yellow arrows represent relative pose edges $(i,j)$. In the OLC space, the object’s center aligns with the origin and has a semantically meaningful orientation. For example, the opening of a cup or bottle aligns with the y-axis, while the front and sides of the object are aligned with the x-axis and z-axis, respectively.
TABLE II: Comparison of COPE methods on the COPE-119 dataset. Methods are categorized into those that predict NOCS maps and those that directly predict object poses. * indicates that only the NOCS map prediction component was used. denotes that the method does not predict scale.
Method IoU25 $\uparrow$ IoU50 $\uparrow$ IoU75 $\uparrow$ 5°2cm $\uparrow$ 5°5cm $\uparrow$ 10°2cm $\uparrow$ 10°5cm $\uparrow$
NOCS [8] 73.50 48.75 20.47 11.57 16.71 22.43 32.98
NOCS (MV-ROPE)* [49] 88.74 71.67 38.19 28.84 31.94 48.50 55.53
DiffusionNOCS [32] 83.00 38.89 4.57 39.87 43.83 55.82 64.80
VI-Net [39] 95.66 87.34 52.35 55.09 59.98 72.35 79.70
IST-Net [35] 97.52 90.68 58.99 51.57 52.85 70.90 73.41
GenPose [37] - - - 47.48 54.30 60.96 71.84
SecondPose [40] 98.55 90.90 53.83 58.57 64.00 75.13 83.29
AGPose [41] 98.47 92.94 66.34 61.61 63.24 80.02 82.06
TABLE III: Comparison of Fine-tuned VS. Original COPE methods on COPE-119 test split, * indicates fine-tuned methods.
Method IoU25 $\uparrow$ IoU50 $\uparrow$ IoU75 $\uparrow$ 5°2cm $\uparrow$ 5°5cm $\uparrow$ 10°2cm $\uparrow$ 10°5cm $\uparrow$
SecondPose [40] 98.53 92.24 57.42 61.88 67.05 78.24 84.08
AGPose [41] 98.52 92.10 68.53 67.11 67.50 81.23 81.92
SecondPose* [40] 99.38(+0.85) 96.23(+3.99) 59.80(+2.38) 76.33(+14.45) 82.91(+15.86) 89.43(+11.19) 97.19(+13.11)
AGPose* [41] 99.91(+1.39) 99.76(+7.66) 84.42(+15.89) 75.10(+8.99) 75.74(+8.24) 94.35(+13.12) 96.46(+14.54)
IV EVALUATION
IV-A Overview
As discussed in  section II-A1, IOPE methods are not our primary focus, instead, we adopt the COPE and UOPE methods for benchmarking, which are capable of generalizing across intra-class variations and unseen objects without the need for individual retraining for each object. These methods can directly leverage open-source pre-trained weights. This approach serves two purposes: first, it allows us to verify the validity and accuracy of our data; second, it helps evaluate the generalization capability of the open-source algorithm on our dataset. Furthermore, we fine-tuned several variants of the COPE method on the COPE-119 training set and evaluated their performance on the test set to validate the effectiveness of our ground truth labels for network training. Additionally, we assessed various object-pose-tracking methods to further broaden our analysis.
IV-B Evaluation Metrics
We follow [8] and adopt the evaluation protocol commonly used for COPE methods, which includes the mean precision of 3D intersection over union (IoU) to jointly assess rotation, translation, and size prediction. 3D IoU measures the overlap between the 3D bounding box of an object derived from the predicted pose and the ground truth bounding box, providing a comprehensive metric. For rotation and translation errors, we typically report the accuracy within a specified threshold. For instance, RotmTransn represents the percentage of predicted poses with a rotation error less than $m$ degrees and a translation error less than $n$ centimeters.
We evaluated the Average Recall (AR) [50] for UOPE methods using three error functions: Visible Surface Discrepancy (VSD), Maximum Symmetry-Aware Surface Distance (MSSD), and Maximum Symmetry-Aware Projection Distance (MSPD). In addition, we report the frame rate (FPS) for all UOPE methods.
The evaluation of OPT methods performance relies on two complementary metrics: We evaluate object pose tracking performance using Absolute Trajectory Error (ATE) for global accuracy and Relative Pose Error (RPE) [51] for local accuracy, reporting average results across all 175 sequences. We also report the AUC of ADD and ADD-S [11], as well as the average recall of ADD(S) within 10% of the object diameter (termed ADD(S)-0.1d) [52].
IV-C Category-level Pose Estimation
IV-C1 COPE-119 Baseline Evaluation
For COPE task, we evaluate 8 methods:  [8],  [49], and  [32] predict canonical representations directly, while  [39],  [35],  [37],  [40], and  [41] predict object pose and category. Except for NOCS [8], which performs joint object detection and pose estimation, all methods utilize identical SAM2 [45] segmentation results to ensure a fair benchmark. Notably, DiffusionNOCS [32] achieves a relatively lower IoU score but exhibits a very small pose error. A possible reason is that it has never been trained on real-world data, leading to instability in object scale prediction.
IV-C2 COPE-119 Fine-Tuning Evaluation
The purpose of this experiment is to demonstrate the accuracy of pose annotations, based on the assumption that the metric improves only when the ground truth pose annotations in both the training and test sets are correct. We fine-tuned AGPose [41] and SecondPose [40] on our 84-sequence training set for 2 epochs and evaluated them on our 35-sequence test set. As shown in  table III, the fine-tuned results significantly outperform their original test set performance by a large margin in merely 2 epochs of fine-tuning, highlighting the accuracy of pose annotations in our dataset.
TABLE IV: Comparison of UOPE methods on the UOPE-56 datasets. * indicates pose estimation without refinement.
Method Input $\uparrow$ VSD $\uparrow$ MSSD $\uparrow$ MSPD $\uparrow$ AR $\uparrow$ FPS $\downarrow$
FoundPose* [30] RGB 50.12 52.02 79.25 60.46 0.50
MegaPose [28] RGB 64.56 68.45 83.64 72.23 2.25
GigaPose [29] RGB 64.84 61.32 83.20 69.79 3.37
MegaPose [28] RGB-D 76.34 80.48 84.64 80.49 2.61
SAM6D [25] RGB-D 79.67 77.95 81.65 79.75 3.75
FoundationPose [1] RGB-D 92.74 90.09 94.40 92.41 0.98
IV-D Unseen Object Pose Estimation
We evaluate 5 UOPE methods in table table IV, grouped by their input modalities into RGB-based and RGB-D-based approaches. To ensure fairness, each sequence is paired with a specific CAD model and uniform segmentations produced by SAM2 [45]. Moreover, methods that rely on CNOS [53] for detection or segmentation also utilize SAM2. Except for SAM6D [25], which experiences performance bottlenecks due to its intrinsic segmentation component and fixed threshold settings that fail to generalize in our dataset’s cluttered backgrounds, we apply dynamic threshold adjustments to maintain accurate pose estimation. MegaPose [28], available in RGB-only and RGB-D configurations, leverages extensive pre-processing to extract rich template information, demonstrating a significant advantage, especially on single-object sequences with ground-truth detection. Notably, FoundPose [30] does not open-source its core refinement code, limiting its evaluation of the results, so we only report its coarse results.
IV-E Object Pose Tracking
For object pose tracking methods, we benchmark  [54],  [42],  [4],  [43], and  [30]. We mainly employed the original algorithms, but with DROID-SLAM [7], we integrated SAM2-derived masks [45] to zero out weights of non-object regions directly within the DBA layer to derive the camera-to-object poses, effectively adapting it for object pose tracking. We report FoundationPose [1] results under the model-based mode.
TABLE V: Comparison on the DynOPETs dataset. The evaluation of average ATE (m), RPE on rotation (Rad), and average RPE on translation (m).
Method ATE $\downarrow$ RPE Rot. $\downarrow$ RPE Trans. $\downarrow$
MaskFusion [54] 0.320 0.060 0.033
DROID-SLAM [4] 0.173 0.063 0.040
BundleTrack [42] 0.068 0.037 0.021
BundleSDF [43] 0.094 0.044 0.026
FoundationPose [1] 0.037 0.018 0.010
TABLE VI: Comparison on DynOPETs dataset. ADD(S) represents AUC (0, 0.1m), and ADD(S)-0.1d is the average recall within 10% of the object diameter.
Method ADD $\uparrow$ ADDS $\uparrow$ ADD-0.1d $\uparrow$ ADDS-0.1d $\uparrow$
MaskFusion [54] 12.23 34.47 28.75 64.65
Droid-SLAM [4] 34.27 57.20 62.20 86.41
BundleTrack [42] 76.39 89.36 94.30 99.54
BundleSDF [43] 70.64 89.41 92.30 99.98
FoundationPose [1] 89.80 96.40 92.42 99.29
By considering both ATE and RPE, we obtain a comprehensive understanding of the pose estimation performance in table table V, encompassing both global consistency and local accuracy. Besides, we benchmark each method’s performance in table table VI across all 175 sequences (COPE-119 & UOPE-56) using the AUC of ADD(S) and ADD(S)-0.1d. Due to failure cases in some sequences, FoundationPose [1] underperforms—especially in ADD(S)-0.1d compared to BundleTrack [42] and BundleSDF [43]. Similarly, MaskFusion [54] shows poor performance, consistent with the trajectory error metrics.
V CONCLUSION
In this work, we propose an efficient framework for acquiring and annotating object pose estimation datasets in environments where both objects and cameras are in motion. Leveraging this approach and our environmental assumptions, we introduce a novel dataset tailored for various object pose tasks. We evaluate 18 state-of-the-art algorithms and provide reference performance benchmarks, validating the high value of the dataset. Future work includes deploying this data acquisition pipeline on a real-world robotic platform to further advance research in robot grasping and manipulation. Another promising direction is to augment the dataset with human hand pose annotations to facilitate research in AR/MR interaction.